# Talking Heads: Understanding Inter-layer Communication in Transformer Language Models

### v2

@@ 1 Introduction
Specifically, by building on recent work in circuit analysis [Elhage et al., 2021, Wang et al., 2022, Goldowsky-Dill et al., 2023, Quirke and Barez, 2024, Merullo et al., 2024, Hanna et al., 2023], we demonstrate how a Transformer LM (GPT2 small, Pythia 160m) passes information from early layers to later ones using low-rank subspaces. These communication channels are proposed in Elhage et al. [2021], but understanding their implementation in weights is an open challenge [Makelov et al., 2024]. This conflicts with circuit analysis, which supports the interpretation of specialization and communication between specific transformer components. Elhage et al. [2021] introduce a score for measuring how much weight matrices in a transformer communicate, but outside of toy models, it has been difficult to interpret the results [Singh et al., 2024]. We build on this work by proposing a method using the Singular Value Decomposition to find these channels by decomposing one matrix into all of its component signals, and find it is much more interpretable than the non-decomposed variant.We focus on two examples of such channels (inhibition and duplicate detection, §2), and find that they are are very low-rank (1 or 2 dimensions), easily interpretable, and causally important for specific model behaviors. We also show this method can be used to perform model editing at training time (Appendix H.2), and provide some encouraging early evidence we can perform circuit discovery without running models (Appendix I). Specifically, our contributions are as follows:
@@ 3 Identifying Communication Channels in Low-Rank Subspaces
In this section, we test the hypothesis that model components like attention heads communicate through signals in low-rank subspaces of the residual stream and that we can find these signals in the weights themselves. We investigate one case of each type of composition outlined in Section 2 and find positive evidence for this hypothesis with query and value composition in 1 and 2 dimensional subspaces, but not key composition (see Appendix C). Because we are able to localize the query and value signals to such small representation spaces, we find that we are able to interpret and control these features with intervention experiments in Section 4.
@@ 3.2 Composition with Decomposed Component Matrices
We initially use the composition score in Equation 1 to try predict these interactions in the weights, but find these results are largely noisy and uninterpretable. This is briefly demonstrated in Appendix I. We find that despite empirically knowing interactions exist between heads, we do not find they reliably have higher composition scores than any random head. Although the composition score has been shown to be useful on small toy models [Elhage et al., 2021], previous work has also shown that on larger models, the signal the composition score conveys is extremely noisy [Singh et al., 2024]. Therefore, we turn to the Singular Value Decomposition (SVD), defined as $\mathbf{W}=\mathbf{USV^{T}}$, on the QK and OV matrices to decompose the attention heads into orthogonal components which determine the input and output spaces of the matrix. This allows us to individually view subspaces read from/written into ordered by the amount of variance of the transformation of the matrix they account for. This helps us answer our original hypothesis that model components communicate across layers in low-rank subspaces of the residual stream.
@@
We find that decomposing weight matrices this way is very effective at cleaning up the composition score signal. We find attention heads that have very high relative composition scores with one component matrix of another head. For example, the second component of head 8.6 (referred to as 8.6.2) composes far more with mover head 9.9 than any of the other 63 components in 8.6 (5 standard deviations higher than the average) or when considering the full matrix as in $CS(\mathbf{OV}^{8.6},\mathbf{QK}^{9.9})$. The bottom graphs in Figure 2 show these results for the inhibition heads. All of the inhibition head exhibit a similar phenomenon of single component dominance. The duplicate token head 3.0 also value-composes with inhibition heads in a similar way, using two components (3.0.1 and 3.0.2) far more than any other. Results are shown in Appendix D.
We can also use this decomposition to find specific pairs of heads that talk more than others. With the knowledge that two heads talk through a specific component of one head, we can find the other heads that communicate through this pathway. Doing so lets us find the signal encoding almost the full IOI circuit in GPT2-Small directly from the weights, without running the model. We outline these results in Appendix I.
@@ 3.3.2 Results
Our editing technique is simply to zero out one component at a time test how this affects copying behavior across a dataset of IOI examples. One way to think about this is zeroing out one singular value of e.g., the OV or matrix, or subtracting one of the component matrices from the sum in Equation 2. We must make the edit to the decomposition and then split the matrices back out so that we can run the model. Given $\mathbf{OV}=\mathbf{USV}^{T}$, after zeroing out some singular value in $\mathbf{S}$ (forming $\mathbf{S^{\prime}}$) we can set the Output and Value matrices to be $\mathbf{U}\sqrt{\mathbf{S^{\prime}}}$ and $\sqrt{\mathbf{S^{\prime}}}\mathbf{V}^{T}$, respectively. In the top graphs of Figure 2, we show that removing the speculated communication channel from the inhibition heads almost always results in a significant decrease in the model’s ability to pass the inhibition signal, with the exception that changing 7.3 does not have a strong effect on its own. In general removing the single component with the highest composition score reduces the Inhibition score by 7-14%, and it is important to consider that this is only when changing a single component in one head at a time. We perform additional experiments with removing/modifying multiple of these components in Section 4 and Appendix E. Thus, we have both behavioral and weight based evidence converging on the interpretation of these subspaces as meaningful communication channels.
@@ 4.1 Interventions on Communication Channels
Figure 4: We find that the 1D inhibition components and 2D duplicate token components finely control which name is avoided by the mover head. On the top, we can selectively inhibit either the first or second name depending on how we scale a vector lying on the 8.10.1 output space. This is strictly controlling relative position. On the bottom, we find that adding or removing duplicate token information from the duplicate channel at the IO or S1 tokens also effectively modulates which name is inhibited. Neither random heads, nor non-communication channel components exhibit these same effects (right). See Appendix E for results on other heads.
@@
Although we get fine control over the attention of the mover head, we have not answered whether this has a real effect on the output behavior. Additionally, Makelov et al. [2024] argue that interventions on subspaces, such as the case here, are prone to a subspace illusion in which the effect is not what it seems. To address this, we measure the Fractional Logit Difference Decrease (FLDD) and Interchange Accuracy of patching the subspaces in the inhibition channel between minimally different IOI examples. GPT2-Small achieves an FLDD of 97.5% and and interchange accuracy of 35%. Makelov et al. [2024] report a baseline that achieves -8% FLDD and 0.0% Interchange. By taking the gradient to directly optimize a single vector for this task they achieve 111.5% FLDD and 45.1% Interchange. Our results in comparison, support the view that these 1D subspaces are a primary mechanism controlling name selection. In Appendix G we also find that these inhibition signals are active broadly during next work prediction. On OpenWebText [Gokaslan and Cohen, 2019], we find that the inhibition heads are primarily active in lists and settings where repetitions should be avoided; for example, in comma separated lists (attending from commas to previously seen items). A natural followup given the IOI results and these observations is to explore their role as an indexing function, which we perform with the Laundry List task.
@@ 5.1 Laundry List Task
We propose a synthetic task that is designed to activate the inhibition heads, but allows us to test their effect on an arbitrary number of candidate tokens. An example is given in Figure 1 and more details on how we generated the data are in Appendix F. Each example first mentions N objects, then N-1, and the model must predict the missing item. We create a dataset of 250 prompts for each value of N from 3 to 20.
@@ 6 Related Work
The SVD has been used for interpretability and weight based analysis in the past [Sule et al., 2023, Praggastis et al., 2022, beren and Black, 2022]. Martin et al. [2021], for example, use the SVD on weight matrices to measure weiht quality to predict generalization performance and whether or not a model is well-trained. The SVD has been used in the training of LoRA modules [Hu et al., 2022], as well as in other finetuning methods [Bałazy et al., 2024, Feng et al., 2024, Wang et al., 2024, Sun et al., 2024, Guo et al., 2024, Karimi Mahabadi et al., 2021]. Our work analyzes model properties that may facilitate these methods to work. Aghajanyan et al. [2021] is perhaps the first to report that pretrained models have a very low intrinsic dimensionality, which helps explain and support our claim that we see so many low rank communication channels in large models. We are therefore excited about the connections between this line of work and work on the linear representations in language models [Elhage et al., 2022, Park et al., 2023], which argues that features in LMs are represented as directions in space. There has been recent work in studying interactions between components in finer-grained feature spaces in the past [Kim et al., 2018, Marks et al., 2024, Geiger et al., 2023, Lepori et al., 2023, Zhang et al., 2024], whereas our approach begins by first analyzing the weights to find substantial connections in subspaces without requiring any data. Future work could connect these two, to find circuits, subnetworks, and/or directions in space for certain behaviors informed by connections in the weights, which alleviates the concern of finding these things from scratch. We outline preliminary results for such an approach for novel circuit discovery in Appendix I.
@@ Appendix A Limitations
Appendix B Larger Models’ Performance When Increasing the # of Objects in the Laundry List Task
@@
Appendix C The Composition Score with and without Weight Decomposition
@@
Appendix D Duplicate Token Heads
@@
Appendix E More IOI Interventions
@@
Appendix F Laundry List Data Generation
@@
Appendix G Inhibition Heads Behavior on Open-domain Text
@@
Appendix H Inhibition Mechanism in Pythia and Training Progression of Inhibition
@@ H.1 Path Patching on IOI
H.2 Training Progression of Inhibition Components
@@
Appendix I Circuit Discovery with Static Weight Analysis
@@
I.1 Finding the IOI Circuit in GPT-2 Weights
@@ I.2 Discovering New Connections
I.2.1 Attention Pattern Analysis of 6.1, 6.2, 6.7, and 8.3
We perform a similar attention pattern analysis as shown in Section G on these heads. We run the model on documents from OpenWebText-10K and look at contexts in which these heads attend strongly (.$\geq.5$ of probability mass) to some token. We find that 6.2 is not interpretable in this way, as its attention patterns are more diffuse, but examples for the other heads are shown in Figures 21, 22, and 20. We find qualitatively similar patterns in these heads to inhibition heads, though with less selection for attending to previous items in lists. Head 8.3 is of particular interest because it has the strongest composition with 9.9, and we will focus on it for the remainder of this section. This head also has some particularly salient motifs in its attention patterns: much of the time it appears to attend to a token (or the last token in a multitoken phrase) in context that could be a plausible continuation of the current text. In the following section we explore what the possible function of this head is, and its connection to the 9.9 mover head.
@@
I.2.2 Attention Head 8.3 as a Relevant Context Head
@@
Appendix J More Information on Composition
J.1 Value Composition: Duplicate Token Heads
@@
J.2 Query Composition: Inhibition Heads
@@
J.3 Key Composition: Induction Heads
@@
Appendix K Extra Laundry List Interventions

### v3

@@ 1 Introduction
Specifically, by building on recent work in circuit analysis [Elhage et al., 2021, Wang et al., 2022, Goldowsky-Dill et al., 2023, Quirke and Barez, 2024, Merullo et al., 2024, Hanna et al., 2023], we demonstrate how a Transformer LM (GPT2 small, Pythia 160m) passes information from early layers to later ones using low-rank subspaces. These communication channels are proposed in Elhage et al. [2021], but understanding their implementation in weights is an open challenge [Makelov et al., 2024]. This conflicts with circuit analysis, which supports the interpretation of specialization and communication between specific transformer components. Elhage et al. [2021] introduce a score for measuring how much weight matrices in a transformer communicate, but outside of toy models, it has been difficult to interpret the results [Singh et al., 2024]. We build on this work by proposing a method using the Singular Value Decomposition to find these channels by decomposing one matrix into all of its component signals, and find it is much more interpretable than the non-decomposed variant.We focus on two examples of such channels (inhibition and duplicate detection, §2), and find that they are are very low-rank (1 or 2 dimensions), easily interpretable, and causally important for specific model behaviors. We also show this method can be used to perform model editing at training time (Appendix G.2), and provide some encouraging early evidence we can perform circuit discovery without running models (Appendix H). Specifically, our contributions are as follows:
@@ 3 Identifying Communication Channels in Low-Rank Subspaces
In this section, we test the hypothesis that model components like attention heads communicate through signals in low-rank subspaces of the residual stream and that we can find these signals in the weights themselves. We investigate one case of each type of composition outlined in Section 2 and find positive evidence for this hypothesis with query and value composition in 1 and 2 dimensional subspaces, but not key composition (see Appendix B). Because we are able to localize the query and value signals to such small representation spaces, we find that we are able to interpret and control these features with intervention experiments in Section 4.
@@ 3.1 Composition Score
$CS(\mathbf{W^{3}_{I}},\mathbf{W_{O}^{1}})=\frac{||\mathbf{W^{3}_{I}W_{O}^{1}}||_{F}}{||\mathbf{W^{3}_{I}}||_{F}*||\mathbf{W_{O}^{1}}||_{F}}$ (2)
@@ 3.2 Composition with Decomposed Component Matrices
We initially use the composition score in Equation 2 to try predict these interactions in the weights, but find these results are largely noisy and uninterpretable. This is briefly demonstrated in Appendix H. We find that despite empirically knowing interactions exist between heads, we do not find they reliably have higher composition scores than any random head. Although the composition score has been shown to be useful on small toy models [Elhage et al., 2021], previous work has also shown that on larger models, the signal the composition score conveys is extremely noisy [Singh et al., 2024]. Therefore, we turn to the Singular Value Decomposition (SVD), defined as $\mathbf{W}=\mathbf{USV^{T}}$, on the QK and OV matrices to decompose the attention heads into orthogonal components which determine the input and output spaces of the matrix. This allows us to individually view subspaces read from/written into ordered by the amount of variance of the transformation of the matrix they account for. This helps us answer our original hypothesis that model components communicate across layers in low-rank subspaces of the residual stream.
@@
We find that decomposing weight matrices this way is very effective at cleaning up the composition score signal. We find attention heads that have very high relative composition scores with one component matrix of another head. For example, the second component of head 8.6 (referred to as 8.6.2) composes far more with mover head 9.9 than any of the other 63 components in 8.6 (5 standard deviations higher than the average) or when considering the full matrix as in $CS(\mathbf{OV}^{8.6},\mathbf{QK}^{9.9})$. The bottom graphs in Figure 2 show these results for the inhibition heads. All of the inhibition head exhibit a similar phenomenon of single component dominance. The duplicate token head 3.0 also value-composes with inhibition heads in a similar way, using two components (3.0.1 and 3.0.2) far more than any other. Results are shown in Appendix C.
We can also use this decomposition to find specific pairs of heads that talk more than others. With the knowledge that two heads talk through a specific component of one head, we can find the other heads that communicate through this pathway. Doing so lets us find the signal encoding almost the full IOI circuit in GPT2-Small directly from the weights, without running the model. We outline these results in Appendix H.
@@ 3.3.2 Results
Our editing technique is simply to zero out one component at a time test how this affects copying behavior across a dataset of IOI examples. One way to think about this is zeroing out one singular value of e.g., the OV or matrix, or subtracting one of the component matrices from the sum in Equation 3. We must make the edit to the decomposition and then split the matrices back out so that we can run the model. Given $\mathbf{OV}=\mathbf{USV}^{T}$, after zeroing out some singular value in $\mathbf{S}$ (forming $\mathbf{S^{\prime}}$) we can set the Output and Value matrices to be $\mathbf{U}\sqrt{\mathbf{S^{\prime}}}$ and $\sqrt{\mathbf{S^{\prime}}}\mathbf{V}^{T}$, respectively. In the top graphs of Figure 2, we show that removing the speculated communication channel from the inhibition heads almost always results in a significant decrease in the model’s ability to pass the inhibition signal, with the exception that changing 7.3 does not have a strong effect on its own. In general removing the single component with the highest composition score reduces the Inhibition score by 7-14%, and it is important to consider that this is only when changing a single component in one head at a time. We perform additional experiments with removing/modifying multiple of these components in Section 4 and Appendix D. Thus, we have both behavioral and weight based evidence converging on the interpretation of these subspaces as meaningful communication channels.
@@ 4.1 Interventions on Communication Channels
Figure 4: We find that the 1D inhibition components and 2D duplicate token components finely control which name is avoided by the mover head. On the top, we can selectively inhibit either the first or second name depending on how we scale a vector lying on the 8.10.1 output space. This is strictly controlling relative position. On the bottom, we find that adding or removing duplicate token information from the duplicate channel at the IO or S1 tokens also effectively modulates which name is inhibited. Neither random heads, nor non-communication channel components exhibit these same effects (right). See Appendix D for results on other heads.
@@
Although we get fine control over the attention of the mover head, we have not answered whether this has a real effect on the output behavior. Additionally, Makelov et al. [2024] argue that interventions on subspaces, such as the case here, are prone to a subspace illusion in which the effect is not what it seems. To address this, we measure the Fractional Logit Difference Decrease (FLDD) and Interchange Accuracy of patching the subspaces in the inhibition channel between minimally different IOI examples. GPT2-Small achieves an FLDD of 97.5% and and interchange accuracy of 35%. Makelov et al. [2024] report a baseline that achieves -8% FLDD and 0.0% Interchange. By taking the gradient to directly optimize a single vector for this task they achieve 111.5% FLDD and 45.1% Interchange. Our results in comparison, support the view that these 1D subspaces are a primary mechanism controlling name selection. In Appendix F we also find that these inhibition signals are active broadly during next work prediction. On OpenWebText [Gokaslan and Cohen, 2019], we find that the inhibition heads are primarily active in lists and settings where repetitions should be avoided; for example, in comma separated lists (attending from commas to previously seen items). A natural followup given the IOI results and these observations is to explore their role as an indexing function, which we perform with the Laundry List task.
@@ 5.1 Laundry List Task
We propose a synthetic task that is designed to activate the inhibition heads, but allows us to test their effect on an arbitrary number of candidate tokens. An example is given in Figure 1 and more details on how we generated the data are in Appendix E. Each example first mentions N objects, then N-1, and the model must predict the missing item. We create a dataset of 250 prompts for each value of N from 3 to 20.
@@ 6 Related Work
The SVD has been used for interpretability and weight based analysis in the past [Sule et al., 2023, Praggastis et al., 2022, beren and Black, 2022]. Martin et al. [2021], for example, use the SVD on weight matrices to measure weiht quality to predict generalization performance and whether or not a model is well-trained. The SVD has been used in the training of LoRA modules [Hu et al., 2022], as well as in other finetuning methods [Bałazy et al., 2024, Feng et al., 2024, Wang et al., 2024, Sun et al., 2024, Guo et al., 2024, Karimi Mahabadi et al., 2021]. Our work analyzes model properties that may facilitate these methods to work. Aghajanyan et al. [2021] is perhaps the first to report that pretrained models have a very low intrinsic dimensionality, which helps explain and support our claim that we see so many low rank communication channels in large models. We are therefore excited about the connections between this line of work and work on the linear representations in language models [Elhage et al., 2022, Park et al., 2023], which argues that features in LMs are represented as directions in space. There has been recent work in studying interactions between components in finer-grained feature spaces in the past [Kim et al., 2018, Marks et al., 2024, Geiger et al., 2023, Lepori et al., 2023, Zhang et al., 2024], whereas our approach begins by first analyzing the weights to find substantial connections in subspaces without requiring any data. Future work could connect these two, to find circuits, subnetworks, and/or directions in space for certain behaviors informed by connections in the weights, which alleviates the concern of finding these things from scratch. We outline preliminary results for such an approach for novel circuit discovery in Appendix H.
@@ 8 Limitations
9 Acknowledgments
Thank you to the members of the LUNAR Lab for valuable feedback on this project. This work was generously supported by the Robert J. & Nancy D. Carney Institute for Brain Science under the T32 grant: Training Program for Interactionist Cognitive Neuroscience (ICoN).
Appendix A Larger Models’ Performance When Increasing the # of Objects in the Laundry List Task
@@
Appendix B The Composition Score with and without Weight Decomposition
@@
Appendix C Duplicate Token Heads
@@
Appendix D More IOI Interventions
@@
Appendix E Laundry List Data Generation
@@
Appendix F Inhibition Heads Behavior on Open-domain Text
@@
Appendix G Inhibition Mechanism in Pythia and Training Progression of Inhibition
@@ G.1 Path Patching on IOI
G.2 Training Progression of Inhibition Components
@@
Appendix H Circuit Discovery with Static Weight Analysis
@@
H.1 Finding the IOI Circuit in GPT-2 Weights
@@ H.2 Discovering New Connections
H.2.1 Attention Pattern Analysis of 6.1, 6.2, 6.7, and 8.3
We perform a similar attention pattern analysis as shown in Section F on these heads. We run the model on documents from OpenWebText-10K and look at contexts in which these heads attend strongly (.$\geq.5$ of probability mass) to some token. We find that 6.2 is not interpretable in this way, as its attention patterns are more diffuse, but examples for the other heads are shown in Figures 21, 22, and 20. We find qualitatively similar patterns in these heads to inhibition heads, though with less selection for attending to previous items in lists. Head 8.3 is of particular interest because it has the strongest composition with 9.9, and we will focus on it for the remainder of this section. This head also has some particularly salient motifs in its attention patterns: much of the time it appears to attend to a token (or the last token in a multitoken phrase) in context that could be a plausible continuation of the current text. In the following section we explore what the possible function of this head is, and its connection to the 9.9 mover head.
@@
H.2.2 Attention Head 8.3 as a Relevant Context Head
@@
Appendix I More Information on Composition
I.1 Value Composition: Duplicate Token Heads
@@
I.2 Query Composition: Inhibition Heads
@@
I.3 Key Composition: Induction Heads
@@
Appendix J Extra Laundry List Interventions